OpenStudy (anonymous):

Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f − g)(3)? 117 101 16 −101

OpenStudy (anonymous):

101 (13(3)^2) - (7(3)-5) = 117-16 = 101

OpenStudy (anonymous):

Consider the functions f(x) = 8x + 2 and g(x) = 5x − 3. What is the value of ( f/g )(4)? Please show all of your work for full credit.

OpenStudy (anonymous):

2 (8(4)+2) / (5(4)-3) = 34/17 = 2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome